The stockholders’ equity section of Concord Corporation’s balance sheet at December 31 is presented here.

CONCORD CORPORATION
Balance Sheet (partial)
Stockholders’ equity
   Paid-in capital
      Preferred stock, cumulative, 11,300 shares authorized, 6,800 shares issued and outstanding

$ 707,200

      Common stock, no par, 739,000 shares authorized, 569,000 shares issued

2,845,000

   Total paid-in capital

3,552,200

   Retained earnings

1,184,000

        Total paid-in capital and retained earnings

4,736,200

   Less: Treasury stock (7,000 common shares)

37,000

Total stockholders’ equity

$4,699,200


From a review of the stockholders’ equity section, answer the following questions.

(a) How many shares of common stock are outstanding?

Common stock outstanding

enter a number of shares shares


(b) Assuming there is a stated value, what is the stated value of the common stock?

The stated value of the common stock

$enter a common stock value per share in dollars  per share


(c) What is the par value of the preferred stock?

The par value of the preferred stock

$enter a preferred stock par value per share in dollars  per share


(d) If the annual dividend on preferred stock is $49,504, what is the dividend rate on preferred stock?

The dividend rate

enter a dividend rate on preferred stock in percentages %


(e) If dividends of $75,700 were in arrears on preferred stock, what would be the balance reported for retained earnings?

The Retained Earnings balance

$enter a Retained Earnings balance in dollars
Answer #1
a
Common stock outstanding 562000 =569000-7000
b
The stated value of the common stock 5 =2845000/569000
c
The par value of the preferred stock 104 =707200/6800
d
The dividend rate 7% =49504/707200
e
The Retained Earnings balance 1184000